OpenStudy (boldgeorge):

Explain the steps you would use to solve the following system of equations with elimination method. Do not solve the system. 1. 5x+2y=5 2. 3x-4y=-23 Thanks in advance!

OpenStudy (anonymous):

you have to solve for one variable first. this can be done by multiplying everything in equation 1 by 2, and then add the equations to each other which will leave you an equation with only x in it

OpenStudy (anonymous):

that should get you started

OpenStudy (jhannybean):

Let's say we wanted to solve for y first,and we wanted to get rid of the x we would multiply both equations by the opposite leading coefficients. 3(5x+2y=5) 5(3x-4y=-23) negate 1 of the two equations. this would help to make x easy to get rid of. -3(5x+2y=5) 5(3x-4y=-23) once you eliminate the x's, you can solve for y using the elimination method. once you get your solution for y, plug it back into either of the two equations to solve for x. Only use one equation to plug it back into though. 5x+2(y)=5 or 3x-4(y)=-23 NOT both.